# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ II.5.3 Valley Splitting Statistics
$E_{\mathrm{VS}}\sim\mathrm{Rice}\left(\nu=2\left|\Delta_{\mathrm{det}}\right|,\sigma^{2}=2\,\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle\right),$ (36)
see Appendix C.2 for details. This result has been obtained previously using the $2k_{0}$-theory in Refs. [19, 18] with a very similar model for the alloy disorder. Here, it has been extended to a more complex case that includes non-trivial resonances and strain. We find that the qualitative result of a Rice distribution for $E_{\mathrm{VS}}$ is unchanged under these extensions, but the shape parameters of the distribution $\nu=2\left|\Delta_{\mathrm{det}}\right|$ and $\sigma^{2}=2\,\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$ are modified to account for the more complex physics, see Eqs. (24)–(25) and (34).
The Rice distribution was found to be in good agreement with fully atomistic tight-binding simulations [18, 32], density functional theory [44] and experimental results from conveyor-mode shuttling tomography of the valley splitting [30].
In the following, we will frequently consider the mean and variance of the Rice distribution given by
@@
with the shape parameters $\nu$ and $\sigma$ specified above. The expression for the mean involves the (monotonically increasing) function
@@
where $I_{n}\left(x\right)$ denotes the modified Bessel functions of first kind. The asymptotics
@@
In order to distinguish between regimes with primarily deterministic or disorder-dominated contributions to the mean valley splitting, we consider the separatrix defined by the condition
@@
$\left.\frac{\nu}{2\sigma}\right|_{\mathrm{separatrix}}\approx 0.3507$
@@
$X_{b}$ mean barrier Ge concentration $0.3$
@@
Tab. 1: Parameter values used in the computations, if not stated otherwise. Band structure parameters (including effective masses) were computed as a function of strain using the empirical pseudopotential model described in Ref. [51] with parameters stated there. The QW thickness is given in units of monolayers $\mathrm{ML}=a_{0}/4$ of the relaxed Si crystal.
@@ III Results
In this section, the model described in Sec. II is employed to compute the valley splitting for different types of engineered heterostructures. We will specifically highlight the effects of shear strain and non-trivial resonances, that go beyond previously reported results. Parameter values used in the computations are given in Tab. 1.
Fig. 3: Impact of interface width on valley splitting. (a) Mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ and shape parameters $\nu$ and $\sigma$ of the Rice distribution in the absence of shear strain $\varepsilon_{x,y}=0$ as a function of the QW interface width, cf. Eq. (10). The interface width is given in units of monolayers $\mathrm{ML}=a_{0}/4\approx 0.1357\,\mathrm{nm}$ of relaxed Si. The gray shaded region indicates the deterministically enhanced regime according to Eq. (39) and the red shaded region shows the [25%, 75%] quantile of the Rice distribution. Deterministic enhancements are observed for sharp interfaces with up to one ML width. (b) Same as (a), but for small shear strain $\varepsilon_{x,y}=0.1\%$. The deterministically enhanced regime is extended to about four MLs. (c) Absolute values of the individual components $\Delta_{\mathrm{det},n}$ contributing to the deterministic valley splitting, see Eq. (27), as a function of the interface width. In the absence of shear strain, the resulting valley slitting is dominated by the $n=0$ and $n=-2$ resonances in the deterministically enhanced regime, whereas the $n=0$ resonance dominates in the disorder-dominated regime. (d) In the case of small shear strain $\varepsilon_{x,y}=0.1\%$ the $n=-1$ component is the dominant contribution over the entire parameter domain. In the simulation, an electric field of $F=10\,\mathrm{mV}/\mathrm{nm}$ was assumed.
@@ III.1 Conventional Heterostructure: Dependence on Quantum Well Interface Width
The results in Fig. 3 show indeed, that the influence of shear strain on sharp interfaces can be significant. Without shear strain, extremely sharp interfaces with a width $\sigma_{u,l}\lesssim a_{0}/4$ of less than one monolayer (ML) are required to achieve a deterministic enhancement of the valley splitting, see Fig. 3 (a). This result is consistent with previous findings based on the $2k_{0}$-theory [19, 20]. If shear strain is applied, the range in which a deterministic enhancement can be expected is considerably increased. In Fig. 3 (b) it is shown that already moderate shear strain of $\varepsilon_{x,y}=0.1\%$ yields deterministic enhancements for rather broad interfaces with a width of up to four MLs. Correspondingly, the valley splitting is increased approximately by a factor of two. In any case, sharp interfaces should be combined with a strong electric field across the QW, in order to enhance the weight of the high-frequency Fourier components by a large overlap of the wave function with the interface.
@@ III.2 Unconventional Heterostructures
$x\left(z\right)=\frac{X_{\mathrm{ww}}}{2}\left(1+\cos\left(qz\right)\right),$ (40)
@@ III.2.1 Uniform Germanium Concentration
For $q=0$, the profile in Eq. (40), reduces to a uniform Ge concentration in the QW. In this configuration, which was first proposed in Ref. [18], the mean valley splitting is enhanced because the random Ge concentration corresponds to a flat (white noise) power spectrum in reciprocal space, contributing even at large wave numbers. The corresponding of the valley splitting enhancement is, however, entirely disorder-dominated. Therefore, numerous spin-valley hotspots must be expected [27].
For a sufficient amount of shear strain, when the deterministic component starts to be dominated by the $n=-1$ resonance, cf. Fig. 3 (d), slight enhancements of the deterministic contribution can be expected. These enhancements are, however, much smaller than the disorder-induced component for typical parameters. This is shown in Fig. 4 (a)–(b), where the valley splittings at $q=0$ are practically unchanged even in the presence of shear strain. Moreover, Fig. 5 (b) shows a regime with enhanced mean valley splitting at very low wave numbers, which is clearly outside of the deterministically enhanced regime (indicated by the separatrix condition (39), dashed line). Finally, the typical epitaxial profile and electron density distribution is shown in Fig. 5 (c) along with the power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi^{2}_{0}\left(z\right)$ and the complex-plane distribution of the intervalley coupling parameter $\Delta$, showing that the valley splitting is indeed disorder-dominated.
Fig. 4: Mean valley splitting in the wiggle well heterostructure as a function of wave number. (a) In the absence of shear strain $\varepsilon_{x,y}=0\%$ deterministic enhancements are observed only at $q=2k_{0}$ (short-period wiggle-well), which increases with increasing amplitude $X_{\mathrm{ww}}$. Away from the resonance, the non-zero valley splittings are disorder-dominated. (b) A small amount of shear strain $\varepsilon_{x,y}=0.1\%$ unlocks a new resonance at $q=2k_{1}$ (long-period wiggle-well). (c) Mean (solid lines) and standard deviation (dashed line) of the valley splitting for fixed amplitude $X_{\mathrm{ww}}=5\%$ and different values of shear strain $\varepsilon_{x,y}$. A constant electric field $F=5\,\mathrm{mV}/\mathrm{nm}$ and biaxial tensile strain, see Eq. (15), was assumed in the simulations.
@@ III.2.2 Short-Period Wiggle-Well
The results shown in Fig. 4 feature a strong resonance at $q=2k_{0}$, which grows with increasing Ge amplitude $X_{\mathrm{ww}}$, but is practically independent of shear strain. This configuration is called the short-period wiggle-well. The effect is based on a direct enhancement of the coupling between the valley states within the same Brillouin zone triggered by the periodicity of the confinement potential. The short-period wiggle-well requires a high-frequency Ge modulation with a very short wave length $\lambda=\pi/k_{0}\approx 2.4\,\mathrm{ML}\approx 0.32\,\mathrm{nm}$. The epitaxial growth of such rapidly modulated structures is out of reach with current technology, which is why the short-period wiggle-well is not relevant for practical applications.
Fig. 5: (a) Mean valley splitting of the long-period wiggle-well with fixed wave number $q=0.32\times 2\pi/a_{0}\approx 2k_{1}$ as a function of shear strain $\varepsilon_{x,y}$ and Ge amplitude $X_{\mathrm{ww}}$. The dashed line is the separatrix defined in Eq. (39) that separates disorder-dominated and deterministically enhanced regimes. For large shear strains, a deterministic enhancement is achieved already for very small Ge amplitudes. The red line indicates the constant shear strain value considered in panel (b). (b) Mean valley splitting as a function of wave number $q$ and amplitude $X_{\mathrm{ww}}$ for fixed shear strain $\varepsilon_{x,y}=0.07\%$. Three domains with valley splitting enhancements can be observed, namely, at very low $q\approx 0$ (neary constant uniform Ge concentration in the QW), near $q\approx k_{1}$ (lower harmonic/ Ge spike, only for sufficiently high $X_{\mathrm{ww}}$) and near $q\approx 2k_{1}$ (long-period wiggle-well). The dashed line is again the separatrix. (c) Left: Epitaxial profile and ground state electron density distribution for the uniform Ge concentration in the well ($q=0$, $X_{\mathrm{ww}}=0.05$) at shear strain $\varepsilon_{x,y}=0.07\%$. The ratio $\nu/\left(2\sigma\right)$ indicates that this configuration falls into the disorder-dominated regime. Right: Power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\psi_{0}^{2}\left(z\right)$ as a function of the wave number $k$. The contribution to the valley splitting is dominated by the $2k_{1}$-resonance and is thus larger than predicted by the $2k_{0}$-theory. The inset shows the statistical distribution of the intervalley coupling parameter in the complex plane in the range of $-0.5\,\mathrm{meV}\leq\mathrm{Re}\left(\Delta\right),\mathrm{Im}\left(\Delta\right)\leq 0.5\,\mathrm{meV}$. (d) Same as (c), but at ($q=k_{1}$, $X_{\mathrm{ww}}=0.17$). The product of the potential and the wave function yields a strong Fourier component at $k=2k_{1}$, despite the potential having only half of the frequency $q=k_{1}$. The configuration has a moderate deterministic enhancement $\nu/\left(2\sigma\right)\approx 0.55$. (e) Same as (c), but for a long-period wiggle-well at ($q=2k_{1}$, $X_{\mathrm{ww}}=0.05$). The $q=2k_{1}$-periodic modulation of the potential provides a significant deterministic enhancement of the valley splitting $\nu/\left(2\sigma\right)\approx 0.86$. A constant electric field $F=5\,\mathrm{mV}/\mathrm{nm}$ and biaxial (tensile) strain was assumed in all simulations.
@@ III.2.3 Long-Period Wiggle-Well
In the presence of shear strain, a second resonance emerges at $q=2k_{1}$, see Fig. 4 (b), where $k_{1}$ has been defined in Eq. (30). The peak grows linearly with increasing amplitude $X_{\mathrm{ww}}$. This resonance is associated with the so-called long-period wiggle-well [34, 31, 33], which enhances the coupling of valley states in neighboring Brillouin zones separated by $2k_{1}$. The corresponding wave length is $\lambda\approx 12.4\,\mathrm{ML}\approx 1.68\,\mathrm{nm}$. Epitaxial growth of such structures is feasible using molecular beam epitaxy.
The dependency on shear strain is illustrated in Fig. 4 (c), showing a linear relation between the mean valley splitting and shear strain. In addition, the plot shows that the standard deviation of the valley splitting (dashed lines) is much smaller than the mean value. This indicates a strong deterministic enhancement of the valley splitting achieved by the long-period wiggle well.
The two-parametric plot in Fig. 5 (a) shows the dependency of the mean valley splitting in a long-period wiggle-well on shear strain and the Ge amplitude. The separatrix (dashed line) indicates that a minimum shear strain of about $\varepsilon_{x,y}\approx 0.02\%$ is required to enter the deterministically enhanced regime even for high Ge amplitudes. For larger shear strains above $\varepsilon_{x,y}\gtrsim 0.08\%$, however, the deterministic enhancement sets in already for tiny Ge amplitudes (but then also with very small mean valley splittings). In Fig. 5 (b), the mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ is shown for fixed shear strain $\varepsilon_{x,y}=0.07\%$ as a function of the wave number $q$ and the Ge oscillation amplitude $X_{\mathrm{ww}}$. The broad resonance around $q\approx 2k_{1}$ falls clearly in the deterministically enhanced regime and is relatively robust against deviations from the ideal wave number. The typical epitaxial profile and the corresponding electron density distribution are shown in Fig. 5 (e) along with the PSD and the complex-plane distribution of the intervalley coupling parameter (inset). The strong spectral contribution at $2k_{1}$ leads to a strong deterministic enhancement of the valley splitting such that the complex normal distribution governing the statistical properties is steered away from the origin. Hence, a significant suppression of spin-valley hotspots can be expected. The long-period wiggle-well is therefore a very promising approach to design qubits with deterministically large valley splittings.
@@ III.2.4 Lower Harmonic Wiggle-Well / Ge-Spike
Finally, we note the emergence of a new resonance in Fig. 4 (b) at approximately $q\approx k_{1}$, i.e., about half the wave number of the long-period wiggle-well. This lower harmonic resonance is suppressed in the absence of shear strain and sets in only at fairly high Ge amplitudes. The resonance is also clearly visible in Fig. 5 (b) for Ge amplitudes above $X_{\mathrm{ww}}\geq 12\%$ and falls clearly within the deterministically enhanced regime. The underlying mechanism becomes clear in Fig. 5 (d), which shows an intricate interplay of the periodic heterostructure potential and the electron ground state wave function. While both of them individually have a frequency spectrum dominated by the $k_{1}$-component, their combination effectively triggers the $2k_{1}$-resonance mechanism that underlies the long-period wiggle-well. From the plot it becomes also clear why a large Ge amplitude is required, since this is crucial to achieve the strong confinement of the electron wave function.
@@
Finally, we remark that the second Ge-peak near the lower interface is not essential for the observed effect. It is simply enforced by the parametrization in Eq. (40). Without the second Ge-peak, the structure has strong similarities with the Ge-spike described in Ref. [35].

### v3

@@ II.5.3 Valley Splitting Statistics
$E_{\mathrm{VS}}\sim\mathrm{Rice}\left(\nu=2\left|\Delta_{\mathrm{det}}\right|,\sigma^{2}=2\,\Gamma\right),$ (35)
see Appendix D.2 for a derivation. This result has been obtained previously using the $2k_{0}$-theory in Refs. [19, 18] with a very similar model for the alloy disorder. Here, it has been extended to a more complex case that includes non-trivial resonances and strain. We find that the qualitative result of a Rice distribution for $E_{\mathrm{VS}}$ is unchanged but the shape parameters of the distribution $\nu=2\left|\Delta_{\mathrm{det}}\right|$ and $\sigma^{2}=2\Gamma=2\,\langle\left|\Delta_{\mathrm{rand}}\right|^{2}\rangle$ are modified to account for the more complex physics, see Eqs. (24) and (33). The Rice distribution was found to be in good agreement with fully atomistic tight-binding simulations [18, 32], density functional theory [47] and experimental results from conveyor-mode shuttling tomography of the valley splitting [30].
In the following, we will frequently consider the mean and the variance of the Rice distribution given by
@@
with the shape parameters $\nu$ and $\sigma$ specified above. The expression for the mean involves the function
@@
where $I_{n}\left(x\right)$ denotes the modified Bessel functions of the first kind. The asymptotics
@@
In order to distinguish between regimes with primarily deterministic or disorder-dominated contributions to the mean valley splitting, we introduce the separatrix defined by the condition
@@
$\left.\frac{\nu}{2\sigma}\right|_{\mathrm{separatrix}}=\left.\frac{\left|\Delta_{\mathrm{det}}\right|}{\sqrt{2\Gamma}}\right|_{\mathrm{separatrix}}\approx 0.3507$
@@
$X_{b}$ mean barrier Ge concentration 0.30.30.3
@@
$C_{1,1}$ longitudinal stiffness constant of Si $167.5\,\mathrm{GPa}$ [54]
$C_{1,2}$ transverse stiffness constant of Si $65.0\,\mathrm{GPa}$ [54]
Tab. 2: Parameter values used in the computations, if not stated otherwise. The QW thickness is given in units of monolayers $\mathrm{ML}=a_{0}/4$ of the relaxed Si crystal.
@@ III Results
In this section, we compute the valley splitting for different types of engineered heterostructures using the model described in Sec. II. We will specifically highlight the effects of shear strain and non-trivial resonances, that go beyond previously reported results. Parameter values used in the simulations are given in Tab. 1 and 2.
Fig. 3: Impact of interface width on valley splitting. (a) Mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ and shape parameters $\nu=2|\Delta_{\mathrm{det}}|$ and $\sigma=\sqrt{2\Gamma}$ of the Rice distribution in the absence of shear strain $\varepsilon_{x,y}=0$ as a function of the QW interface width. The interface width is given in units of monolayers $\mathrm{ML}=a_{0}/4\approx 0.1357\,\mathrm{nm}$ of relaxed Si. The gray shaded region indicates the deterministically enhanced regime according to Eq. (38) and the red shaded region shows the [25%, 75%] quantile of the Rice distribution. Deterministic enhancements are observed for sharp interfaces with up to one ML width. (b) Same as (a), but for small shear strain $\varepsilon_{x,y}=0.05\%$. The deterministically enhanced regime is extended to about five MLs. (c) Absolute values of the individual components $\Delta_{\mathrm{det},n}$ contributing to the deterministic valley splitting, see Eq. (26), as a function of the interface width. In the absence of shear strain, the resulting valley slitting is dominated by the $n=0$ resonance with a small correction induced by the $n=-2$ resonance. (d) In the case of small shear strain $\varepsilon_{x,y}=0.05\%$, the $n=-1$ resonance provides the dominant contribution over almost the entire parameter domain. In the simulation, an electric field of $F=10\,\mathrm{mV}/\mathrm{nm}$ was assumed.
@@ III.1 Conventional Heterostructure: Dependence on Quantum Well Interface Width
The results in Fig. 3 show that the influence of shear strain on sharp interfaces is indeed significant. Without shear strain, extremely sharp interfaces with a width $\sigma_{u,l}\approx a_{0}/4$ of about one monolayer (ML) are required to achieve a deterministic enhancement of the valley splitting, see Fig. 3 (a). This result is consistent with previous findings based on the $2k_{0}$-theory [19, 20]. If shear strain is applied, the range in which a deterministic enhancement occurs is significantly increased. In Fig. 3 (b) it is shown that already moderate shear strain of $\varepsilon_{x,y}=0.05\%$ yields deterministic enhancements for rather broad interfaces with a width of up to five MLs. In any case, sharp interfaces must be combined with a strong vertical electric field to enhance the overlap of the wave function with the interface.
@@ III.2 Unconventional Heterostructures
$x\left(z\right)=\frac{1}{2}X_{\mathrm{ww}}\left(1+\cos\left(qz\right)\right)$ (39)
@@
Fig. 4: Mean valley splitting in the wiggle well heterostructure as a function of wave number. (a) In the absence of shear strain $\varepsilon_{x,y}=0\%$ deterministic enhancements are observed only at $q=2k_{0}$ (short-period wiggle-well), which increases with increasing amplitude $X_{\mathrm{ww}}$. Away from the resonance, the non-zero valley splittings are disorder-dominated. (b) A small amount of shear strain $\varepsilon_{x,y}=0.1\,\%$ unlocks a new resonance at $q=2k_{1}$ (long-period wiggle-well). (c) Mean valley splitting for fixed amplitude $X_{\mathrm{ww}}=5\%$ and different values of shear strain $\varepsilon_{x,y}$. A constant electric field $F=5\,\mathrm{mV}/\mathrm{nm}$ and biaxial tensile strain, see Eq. (15), was assumed in the simulations.
@@ III.2.1 Uniform Germanium Concentration
For $q=0$, the profile in Eq. (39), reduces to a uniform Ge concentration in the QW. In this configuration, which was first proposed in Ref. [18], the mean valley splitting is enhanced because the random Ge concentration corresponds to a flat (white noise) power spectrum in reciprocal space, contributing even at large wave numbers. The corresponding enhancement of the valley splitting is, however, entirely disorder-dominated. Therefore, numerous spin-valley hotspots must be expected [27].
For a sufficient amount of shear strain, when the deterministic component starts to be dominated by the $n=-1$ resonance, cf. Fig. 3 (d), slight enhancements of the deterministic contribution can be expected. These enhancements are, however, much smaller than the disorder-induced component for typical parameters. This is shown in Fig. 4 (a)–(b), where the valley splittings at $q=0$ are practically unchanged even in the presence of shear strain. Moreover, Fig. 5 (b) shows a regime with enhanced mean valley splitting at very low wave numbers, which is clearly outside of the deterministically enhanced regime (indicated by the separatrix condition (38), dashed line). Finally, the typical epitaxial profile and electron density distribution is shown in Fig. 5 (c) along with the power spectral density (PSD) of the product $\left(U_{\mathrm{QW}}\left(z\right)+U_{F}\left(z\right)\right)\left|\psi_{0}\left(z\right)\right|^{2}$ and the complex-plane distribution of the intervalley coupling parameter $\Delta$, showing that the valley splitting is indeed disorder-dominated.
[…]
@@ III.2.2 Short-Period Wiggle-Well
The results shown in Fig. 4 feature a strong resonance at $q=2k_{0}$, which grows with increasing Ge amplitude $X_{\mathrm{ww}}$, but is practically independent of shear strain. This configuration is called the short-period wiggle-well. The effect is based on a direct enhancement of the coupling between the valley states within the same Brillouin zone triggered by the periodicity of the confinement potential. The short-period wiggle-well requires a high-frequency Ge modulation with a very short wave length $\lambda=\pi/k_{0}\approx 2.4\,\mathrm{ML}\approx 0.32\,\mathrm{nm}$. The epitaxial growth of such rapidly modulated structures is out of reach with current growth technology.
@@ III.2.3 Long-Period Wiggle-Well
In the presence of shear strain, a second resonance emerges at $q=2k_{1}$, see Fig. 4 (b). The peak grows almost linearly with increasing amplitude $X_{\mathrm{ww}}$. This resonance is associated with the long-period wiggle-well [34, 31, 33], which enhances the coupling of valley states in neighboring Brillouin zones separated by $2k_{1}$. The corresponding wave length is $\lambda\approx 11.8\,\mathrm{ML}\approx 1.60\,\mathrm{nm}$ (for a biaxially strained QW). Such structures can be grown using molecular beam epitaxy [36].
The dependency on shear strain is illustrated in Fig. 4 (c), showing a linear relation between the mean valley splitting and shear strain at $q=2k_{1}$. The two-parametric plot in Fig. 5 (a) shows the dependency of the mean valley splitting in a long-period wiggle-well on shear strain and the Ge amplitude. The separatrix (dashed line) indicates that a minimum shear strain of about $\varepsilon_{x,y}\approx 0.006\%$ is required to enter the deterministically enhanced regime even for high Ge amplitudes (where disorder effects are strong). For shear strain above $\varepsilon_{x,y}\gtrsim 0.035\%$, the deterministic enhancement sets in already for tiny Ge amplitudes. In Fig. 5 (b), the mean valley splitting $\langle E_{\mathrm{VS}}\rangle$ is shown for fixed shear strain $\varepsilon_{x,y}=0.025\%$ as a function of the wave number $q$ and the Ge oscillation amplitude $X_{\mathrm{ww}}$. The broad resonance around $q\approx 2k_{1}$ falls clearly in the deterministically enhanced regime and is relatively robust against deviations from the ideal wave number. The typical epitaxial profile and the corresponding electron density distribution are shown in Fig. 5 (e) along with the PSD and the complex-plane distribution of the intervalley coupling parameter (inset). The strong spectral contribution at $2k_{1}$ leads to a considerable deterministic enhancement of the valley splitting such that the complex normal distribution governing the statistical properties is steered away from the origin. Hence, a significant suppression of spin-valley hotspots can be expected. The long-period wiggle-well is thus a very promising approach to design qubits with deterministically large valley splittings.
@@ III.2.4 Lower Harmonic Wiggle-Well / Ge-Spike
Finally, we note the emergence of a new resonance in Fig. 4 (b) at approximately $q\approx k_{1}$, i.e., about half the wave number of the long-period wiggle-well. This lower harmonic resonance is suppressed in the absence of shear strain and sets in only at fairly high Ge amplitudes. The resonance is also clearly visible in Fig. 5 (b) for Ge amplitudes above $X_{\mathrm{ww}}\gtrsim 12\%$ and falls clearly within the deterministically enhanced regime. The underlying mechanism becomes clear in Fig. 5 (d), which shows an intricate interplay of the periodic heterostructure potential and the electron ground state wave function. While both of them individually have a frequency spectrum dominated by the $k_{1}$-component, their combination effectively triggers the $2k_{1}$-resonance mechanism that underlies the long-period wiggle-well. From the plot it becomes also clear why a large Ge amplitude is required, since this is crucial to achieve the strong electronic confinement.
@@
Finally, we remark that the second Ge-peak near the lower interface is not essential for the observed effect. It is simply enforced by the parametrization in Eq. (39). Without the second Ge-peak, the structure has similarities with the Ge-spike described in Ref. [35].